OpenStudy (anonymous):

Factor completely: 16x2 − 81

OpenStudy (anonymous):

ANSWERS:

OpenStudy (anonymous):

16x^2 - 81 = 0 (4x + 9)(4x – 9) = 0 So here you look for the solutions. 4x + 9 = 0 and 4x – 9 = 0 x = 9/4 and -9/4

OpenStudy (anonymous):

In other words, the first option is the right answer.